### v1

# AssemAI: Interpretable Image-Based Anomaly Detection for Manufacturing Pipelines

Abstract
Anomaly detection in manufacturing pipelines remains a critical challenge, intensified by the complexity and variability of industrial environments. This paper introduces AssemAI, an interpretable image-based anomaly detection system tailored for smart manufacturing pipelines. Our primary contributions include the creation of a tailored image dataset and the development of a custom object detection model, YOLO-FF, designed explicitly for anomaly detection in manufacturing assembly environments. Utilizing the preprocessed image dataset derived from an industry-focused rocket assembly pipeline, we address the challenge of imbalanced image data and demonstrate the importance of image-based methods in anomaly detection. The proposed approach leverages domain knowledge in data preparation, model development and reasoning. We compare our method against several baselines, including simple CNN and custom Visual Transformer (ViT) models, showcasing the effectiveness of our custom data preparation and pretrained CNN integration. Additionally, we incorporate explainability techniques at both user and model levels, utilizing ontology for user-friendly explanations and SCORE-CAM for in-depth feature and model analysis. Finally, the model was also deployed in a real-time setting. Our results include ablation studies on the baselines, providing a comprehensive evaluation of the proposed system. This work highlights the broader impact of advanced image-based anomaly detection in enhancing the reliability and efficiency of smart manufacturing processes.
Index Terms:
Anomaly Detection, Object Detection, Smart Manufacturing, Interpretability, Image Processing
I Introduction
The evolution of manufacturing has been driven by distinct technological milestones. Initially, mechanization marked the first industrial revolution, followed by the mass production techniques of the second revolution. The third revolution introduced automation and computerization, significantly enhancing operational efficiency. The modern smart manufacturing paradigm emphasizes the utilization of data and advanced analytics to inform decision-making, thereby optimizing productivity and efficiency [1]. The incorporation of the Internet of Things (IoT), artificial intelligence (AI) and other advanced technologies plays a pivotal role in this transformation, revolutionizing manufacturing processes and systems [2].
Anomaly detection is an essential methodology in manufacturing to identify deviations from normal production processes, which can indicate potential issues such as equipment failures, defects, or inefficiencies. Anomaly detection helps ensure product quality, reduce downtime and optimize production processes [3] [4]. By monitoring various process parameters and product characteristics, manufacturers can detect anomalies early and prevent costly production disruptions [5].
Image data is increasingly being leveraged in manufacturing systems due to advancements in computer vision and the availability of high-resolution cameras. In modern manufacturing facilities, image data is utilized for a wide range of applications, including quality inspection [6, 7], predictive maintenance [8, 9], process monitoring [10] and process optimization [11]. The integration of image data allows for real-time monitoring and anomaly detection, significantly improving the ability to identify defects and streamline production processes. Specifically, image-based anomaly detection has gained prominence, leveraging visual data to identify defects or irregularities in products. Automated visual inspection systems use image data to detect surface defects, assembly errors and deviations from design specifications[12], providing real-time feedback essential for maintaining high production standards and minimizing defects[4].
Despite the advancements in image-based anomaly detection, several challenges persist. (i) High variability in appearance between normal and anomalous conditions, which can make it difficult for models to differentiate subtle defects from acceptable variations. (ii) Requirement for large and diverse datasets for training, which can be resource-intensive to collect and annotate [5, 13]. (iii) Challenges with generalization, where a model trained on images from one manufacturing environment may not perform well in different settings due to variations in lighting, camera angles, or product types [14]. (iv) Difficulty in interpreting detection results, affecting the ability to provide actionable insights for process improvement [15].
Figure 1: Overall Architecture. The figure includes the pipeline for AssemAI, beginning with dataset preparation involving filtering and cropping images, followed by zero-shot object detection using two approaches (a custom model and YOLO-FF), and culminating in object detection modeling with CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET. The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations.
To address these challenges, this work presents the following contributions:
  • A novel anomaly detection model utilizing the pre-trained EfficientNet architecture [16], fine-tuned to a new dataset for improved anomaly detection in assembly processes.
  • A new image dataset for anomaly detection in manufacturing, curated using a custom-trained YOLOv9 model named YOLO-FF (You Only Look Once-Future Factories).
  • Integration of advanced explainability techniques, incorporating an ontology-based method for user-level explanations and SCORE-CAM for detailed feature importance and model analysis.
  • The image dataset and code to reproduce the results and additional experiments are available at this link.
II Literature Review
II-A Object Detection Models
Object detection has evolved with various models, improving both accuracy and efficiency. Girshick et al. [17] introduced R-CNN, which applies CNNs to region proposals for object classification and bounding box refinement. Faster R-CNN [18] built on R-CNN by integrating a Region Proposal Network (RPN) for faster detection. The YOLO models, from YOLOv1 to YOLOv9, transformed object detection by formulating it as a single regression problem to predict bounding boxes and class probabilities directly [19]. ShuffleNet [20] and MobileNet [21] are lightweight models designed for efficiency on mobile devices. SqueezeNet [22] aimed to reduce model size while maintaining accuracy with its Fire module. Lastly, the Swin Transformer [23] introduced hierarchical transformers with shifted windows, setting new benchmarks for object detection and vision tasks. Despite these advancements, limitations such as the need for extensive labeled data and computational resources persist. Our work addresses these limitations by focusing on lightweight and efficient models tailored for anomaly detection in manufacturing settings, improving both detection speed and accuracy while maintaining interpretability.
II-B Zero-Shot Object Detection
Zero-shot object detection (ZSD) recognizes objects without labeled training data but faces challenges like semantic noise and class imbalance. Foundational methods by Bansal et al. [24] and Rahman et al. [25] struggled with noise. Gupta et al. [26] and Zheng et al. [27] improved ZSD with symmetric mapping and cascade stages, but issues persisted. Hayat et al. [28] and Li et al. [29] used ResNet, KNN, and contextual information but faced noise and ambiguity. Zhu et al. [30] and Hayat et al. [31] advanced ZSD with DELO and Generative-ZSD. Liu et al. [32] and Li et al. [33] proposed contrastive learning and a semantics-aware framework. Our work focuses on specialized domains for precise anomaly detection.
II-C Anomaly detection
Anomaly detection has seen significant advancements across diverse domains like networking [34], smart agriculture [35], healthcare [36], foundational [35, 37, 38, 39], manufacturing [40, 41, 42, 43, 44, 45]. Kim et al. introduced NetViewer [34], a network measurement approach that detects, identifies and visualizes attacks and anomalous traffic in real-time by passively monitoring packet headers and representing them as frames or images. An anomaly detection method using Generative Adversarial Networks (GAN), which searches for good representations of samples in the generator’s latent space and deems samples without such representations as anomalous, was proposed by Deecke et al. [37]. Minhas et al, in their work addresses visual defect assessment for anomaly detection in tasks [38], like surface inspection by proposing a network-based deep transfer learning method using CNNs, hypothesizing that deeper networks will outperform traditional single class SVMs.
II-D Anomaly detection in manufacturing
Significant research has investigated image-based anomaly detection in manufacturing. Haselmann et al. [39] present an unsupervised one-class learning method using a deep CNN for surface inspection, outperforming other methods on decorated plastic parts. Xie et al. [40] propose a uniform benchmark for assessing image anomaly detection (IAD) algorithms in industrial settings. Maggipinto et al. [41] use convolutional autoencoders for monitoring semiconductor manufacturing, enhancing effectiveness and scalability. Jiang et al. [42] introduce YOLOv3 for balanced datasets and Fast-AnoGAN for unbalanced datasets in industrial production. Tan et al. [43] use an encoder-decoder for anomaly detection in sequential sensor data. Kim et al. [44] propose a self-supervised method using Gramian angular field and StyleGAN for time-series data. Bougaham et al. [45] demonstrate a three-step deep learning approach for Printed Circuit Board Assembly (PCBA) images, achieving high accuracy. Despite advancements, publicly available datasets for assembly processes are scarce, and existing methods lack interpretability and explainability for domain-specific insights.
III Proposed Methodology
Figure 1 shows the overall architecture of the proposed method. The AssemAI pipeline begins with dataset preparation, where images are filtered and cropped to focus on relevant features. This is followed by zero-shot object detection using two approaches: a custom model and YOLO-FF, which adapts zero-shot detection for anomaly detecti use cases. Subsequently, object detection modeling is performed using several architectures, including CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET, to ensure robust anomaly detection. The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations, providing both technical and domain-specific insights.
III-A Future Factories Multimodal Dataset
We use the Future Factories (FF) dataset [46] generated and made publicly available by the Future Factories (FF) lab operating at the McNair Aerospace Research Center at the University of South Carolina. The dataset consists of measurements from a simulation of a rocket assembly pipeline, which adheres to industrial standards in deploying actuators, control mechanisms and transducers. The dataset is comprised of two versions; Analog and Multi-modal dataset, for which we use the images included in the Multi-modal dataset for our purpose. Table I shows the statistics of the original images in the FF Multimodal dataset.
TABLE I: Artificats of the original images in FF Multimodal dataset
Dataset Artifact Statistic
Rarity Percentage 35.73%
Frequency 0.367 Hz
Data collection period 30 hours
Total image count 332002
Original image size 720px*1080px
Types of anomalies 7
Types of classes 8
TABLE II: Artificats of the filtered images in FF Multimodal dataset
Dataset Artifact Statistic
Total image count 15594
Train image count 12475
Test image count 3119
Filtered image
sizes
1:200px*70px (cycle state 4)
2:400px*205px (cycle state 9)
TABLE III: Anomaly types in filtered Multimodal FF dataset
{tblr}
hlines, vlines, Anomaly type & Train
image
count
Test
image
count
Total count
by anomaly
Percentage
No Anomaly 8006 2016 10022 64.26%
NoNose 872 238 1110 7.1%
NoNose,
NoBody2 1222 308 1530 9.8%
NoNose,
NoBody2,
NoBody1 1310 310 1620 10.38%
NoBody1 1065 247 1312 8.4%
Total image
count 12475 3119 15594
III-B Image filtering
The rocket assembly process at the FF lab is divided into 21 distinct cycle states. Information about these cell cycle states was not directly available in the multi-modal dataset and had to be extracted from the analog dataset using a mapping function provided by domain experts. By calculating the Structural Similarity Index (SSIM) between the normal and anomalous images as shown in Figure 2, we observed that the images are structurally very similar across most cycle states. This indicates that the differences between the images are subtle and localized, which suggests that further analysis is needed to filter and figure out the region of interest. Among these cell cycle states, the rocket and its parts are visible only in two specific cycle states due to the spatial location of robots and other machinery and the location of the cameras and camera angles. We focus on filtering cycle four and a section of cycle nine based on domain knowledge and observational insights for enhanced image understanding. Table II and III show the filtered dataset and its anomaly statistics, respectively.
Figure 2: Structural Similarity between Normal and Anomalous Images
III-C Image Cropping
Our dataset comprises images captured at various stages of the manufacturing cycle. Each cycle consists of defined time stamps where the toy rocket’s position remains consistent. Leveraging this domain knowledge, we have implemented a cropping strategy to remove the background from all images. This consistent positioning of the toy rockets across cycles enables us to crop the images accurately, ensuring that only the relevant parts of the images are retained.
The necessity for this cropping approach became evident after our initial attempts at anomaly detection without cropping yielded sub-optimal results. Specifically, our models performed poorly, with significant misclassification rates. We employed the explainability model SCORE-CAM [47] to understand the underlying issues. The insights provided by this model revealed that the models disproportionately focused on the images’ background rather than on the toy rockets themselves, as shown in Figure 3. This misdirection was the primary cause of the poor performance. By cropping the images to remove the background, we ensured that the models concentrated on the relevant features of the toy rockets. This adjustment improved the model performance, leading to more accurate anomaly detection. The final cropped images are of two sizes, determined by the cycle states: 200 pixels by 70 pixels for cycle state 4 and 400 pixels by 205 pixels for cycle state 9.
Figure 3: Score-CAM visualization highlighting the model’s focus on background elements emphasizing the need for cropping and object detection to improve accuracy by isolating the toy rocket.
III-D Text Guided Zero-Shot Object Detection
We then incorporated a zero-shot object detection model to enhance the accuracy and robustness of our anomaly detection pipeline targeting the detection of toy rockets. Specifically, we employed OWL-ViT [48], a model that leverages multi-modal representations to perform open-vocabulary detection. This approach allows for the detection of objects based on free-text prompts, facilitating a flexible and powerful detection mechanism. OWL-ViT integrates CLIP (Contrastive Language-Image Pretraining) with lightweight object classification and localization heads. This integration enables the model to handle open-vocabulary detection by embedding free-text queries through CLIP’s text encoder, which are then utilized as inputs for the object classification and localization heads. The Vision Transformer (ViT) processes image patches as inputs, associating them with their corresponding textual descriptions.
III-E Training a Custom Object Detection Model with YOLOv9
To further enhance our anomaly detection pipeline for toy rockets, we trained a custom object detection model using YOLOv9 [19]. While zero-shot models like OWL-ViT offer significant flexibility and adaptability, a custom-trained model provides specific advantages such as:
  • By training YOLOv9 specifically on our dataset of toy rockets, we can fine-tune the model parameters to achieve higher accuracy and precision for this specific task, compared to a more general-purpose zero-shot model.
  • YOLOv9 is designed for high-speed inference, making it ideal for real-time applications in a manufacturing pipeline. In contrast, zero-shot models like OWL-ViT, while flexible, may have higher computational requirements and longer inference times.
IV Problem Formulation
Consider a dataset $\mathcal{D}$ comprising images captured at various stages of the toy rocket manufacturing process. The manufacturing process is divided into cycles, each consisting of 21 distinct states. Let $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$ denote the set of manufacturing cycles, where each cycle $C_{j}$ for $j=1,2,\ldots,n$ represents the complete assembly of one toy rocket.
Each cycle $C_{j}$ is divided into 21 states, represented by the set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{21}\}$. For our anomaly detection task, we focus on images corresponding to cycle states $S_{4}$ and $S_{9}$, where rocket parts are most likely to be visible.
Let =$\mathcal{I}=\{I_{j,s}\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$ denote the set of images captured during these specific cycle states. Each image $I_{j,s}$ is associated with a label $y_{j,s}\in\mathcal{L}$, where =$\mathcal{L}=\{\text{normal},\text{anomaly}_{1},\text{anomaly}_{2},\ldots,\text{anomaly}_{k}\}$. Additionally, let $B_{j,s}=(x_{\min},y_{\min},x_{\max},y_{\max})$ denote the bounding box for image $I_{j,s}$, obtained via object detection techniques to focus on the relevant part of the image.
We define the dataset as:
$\mathcal{D}=\{(I_{j,s},y_{j,s},B_{j,s})\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$
IV-A Task Description
The task is to develop a model $f:\mathcal{I}\to\mathcal{L}$ that can accurately classify each image $I_{j,s}$ into one of the predefined categories in $\mathcal{L}$. The classification involves the following steps:
  1. 1.
    Filtering the dataset to include only images from cycle states $S_{4}$ and $S_{9}$.
  2. 2.
    Cropping each image $I_{j,s}$ using its bounding box $B_{j,s}$ which is mapped to its specific state. This effectively isolates the regions of interest while removing background noise.
  3. 3.
    Applying AssemAI to classify the cropped images.
The primary objective is to detect anomalies in the images, leveraging the bounding boxes to enhance the model’s ability to focus on the critical parts of the image. To address class imbalance, we use a weighted cross-entropy loss, where higher weights are assigned to classes with fewer images. Mathematically, we aim to minimize the weighted classification error defined as:
$\min_{\theta}\frac{1}{n}\sum_{j=1}^{n}\sum_{s\in\{4,9\}}w_{y_{j,s}}\mathcal{L}(f(I_{j,s};\theta),y_{j,s})$
where $\theta$ represents the model parameters, $\mathcal{L}$ is the cross-entropy loss function and $w_{y_{j,s}}$ is the weight assigned to the class $y_{j,s}$.
By focusing on specific cycle states and using object detection to preprocess the images, we aim to improve the accuracy and reliability of our anomaly detection system. The ultimate goal is to create a robust model that accurately detects anomalies in the toy rocket manufacturing process.
V Experiments
We outline the following experimental setup to evaluate the AssemAI’s overall performance and the contribution of each sub-module.
V-A Common Hyperparameters and Training Setup
The preprocessed dataset was split into training (80%) and testing (20%) sets. For all models, we used Cross-Entropy Loss and the Adam optimizer [49], tuning hyperparameters like epochs, batch size and learning rate. The best model was saved based on validation accuracy. The dataset preprocessing involved resizing images to the required input dimensions and normalizing them according to the specific requirements of each model.
V-B Baselines
V-B1 Custom CNN
The Simple CNN model architecture included two convolutional layers (32 and 64 filters respectively), each followed by a max-pooling layer. After flattening, the output was passed through a fully connected layer with 512 neurons, then to the final layer corresponding to the number of classes (5).
V-B2 Custom ViT
The Vision Transformer (ViT) model [50] was implemented from scratch, starting with a patch embedding layer that splits the input image into non-overlapping patches, each embedded into a high-dimensional space. This was followed by a series of transformer blocks for learning spatial relationships within the image patches. Each transformer block consisted of a multi-head self-attention mechanism and a feed-forward neural network, with layer normalization and residual connections applied at each step. The final classification head mapped the output to the desired number of classes (5).
V-B3 Pretrained ViT
We utilized the pre-trained Vision Transformer (ViT) model, specifically google/vit-base-patch16-224. Images were pre-processed by dividing each image into a sequence of fixed-size non-overlapping patches, then linearly embedded. A [CLS] token was added to represent the entire image, facilitating classification. Absolute position embeddings were incorporated and the resulting sequence of vectors was fed to the standard Transformer encoder. The ViTImageProcessor resized and normalized images to the required 224x224 resolution. The model was loaded with half-precision (torch.float16). Training utilized PyTorch Lightning’s Trainer, with a weighted Cross-Entropy loss function and the AdamW optimizer with a learning rate scheduler. The best checkpoint was saved by monitoring the validation loss.
V-C Proposed Approach
We implemented an EfficientNet-B0 model [16] for image classification tasks. The EfficientNet-B0 architecture was selected for its state-of-the-art performance and efficiency in image classification. It leverages a compound scaling method to optimize model depth, width and resolution, achieving high accuracy with fewer parameters compared to traditional networks. The model was pre-trained on ImageNet and adapted for our task by modifying the final classification layer to match the number of classes in our dataset (5 classes). The model was trained for 10 epochs, with the best model saved based on validation accuracy.
VI Results
Table IV summarizes the results of our experiments on the test set and the ablation studies across different baselines. We evaluate the performance using four metrics: weighted averages of precision, recall, F1-score and accuracy. The weighted averages are calculated based on the four types of anomaly classes and the normal class. The EfficientNet model achieves 96% overall accuracy and 95% of precision, 96% of recall, 95% of F1 score, respectively. Figure 4 depicts the performance of detecting various anomaly types and the normal class across various modeling approaches. It can be observed that among all the models, EfficientNet with the YOLO-FF module gives the best results in detecting all the types of classes.
TABLE IV: Experimental Results
OD Model
WP
WR
WF1
Accuracy Support*
Simple CNN
+YOLO-FF
92.00±
1.00%
91.00±
1.00%
91.00±
1.00%
91.00±
1.00%
3119
Custom ViT
+YOLO-FF
71.00±
1.00%
66.00±
1.00%
66.00±
1.00%
65.00±
1.00%
3119
Pretrained ViT
+YOLO-FF
88.50±
0.50%
87.50±
0.50%
87.50±
0.50%
88.50±
0.50%
3119
EfficientNet with
original images
62.50±
0.50%
60.50±
0.50%
61.50±
0.50%
61.50±
0.50%
332002
EfficientNet with
filtered images
70.50±
0.50%
72.50±
0.50%
73.50±
0.50%
72.50±
0.50%
3119
EfficientNet
+YOLO-FF
95.00±
1.00%
96.00±
1.00%
95.00±
1.00%
96.00±
1.00%
3119
*Support-Number of samples used. WP-Weighted Average Precision, WR-Weighted Average Recall, WF1-Weighted Average F1-Score
Figure 4: Experimental results. Class 1-5 denotes anomaly types: [No Anomaly], [NoNose], [NoNose,NoBody2], [NoNose,NoBody2,NoBody1] and [NoBody1] respectively
VI-A Additional experiments on Explainability
VI-A1 User level explainability
In this work, we employ process ontology designed and developed for Future Factories Rocket assembly line. In comparison with conventional ontologies, process ontology not only captures the definition of sensors but also captures the procedural nature of the assembly process. This aids in understanding the involvement of sensors and equipments at a given point in assembly process. The Future Factories assembly is divided into 21 cycle states which forms the basis of process ontology construction. The specific features of process ontology are as follows: (i) consists of definition and item specification of sensors and equipments, (ii) relationship between the sensors and equipments, (iii) function and involvement of each sensors and robots with respect to the cycle states (iv) expected (or anomalous) values of sensor variables in with respect to each cycle state (v) type of anomaly that could be associated with each cycle state (iv) sensor values and other knowledge can be dynamically updated as per the change in experiment set up. Capturing the procedural nature of the assembly process aids in understanding the contribution of sensors in anomalies. For example, if the initial stage of assembly line is being analyzed for anomaly, it can be understood from ontology that Robot-4 and its corresponding sensors do not contribute to this anomaly.
The goal of process ontology is to explain and assess the output from the models. Given an image, if the model detects an anomaly, the expected values of the sensors present in that image can be obtained and provided to the user. On the other hand, process ontology can also verify if the model predictions are correct in certain cases (Figure 5). Since each image is associated with a timestamp which in turn can be mapped to cycle state, the predicted output from the model can be verified using anomaly types defined in the process ontology. Illustrating using an output from the model using Figure 5, we notice that the input image is associated with cycle state 4. The model predicted the image to be anomalous with type NoNose. This is an incorrect prediction as per the ontology as NoNose anomaly can happen only from cycle state 8 onwards. Using this knowledge and verifying the outputs model predictions, it is found that the model incorrectly predicted NoBody2 anomaly 105 out of 801 times and NoNose anomaly 245 out of 1145 times.
Figure 5: Verification of the explanations through process ontology
VII Deployment of AssemAI
After training and testing the model, we then worked on deploying AssemAI on the Future Factories lab. This deployment strategy can be seen in 6. AssemAI requires two separate types of inputs to operate in real-time. The script that utilizes the model should be able to obtain the current cycle state of the assembly process. This data tag allows the script to ensure that the images captured and input into the model are from cycle states 4 and 9. This tag is obtained by connecting to the OPC UA server running on the Programmable Logic Controller (PLC) and constantly reading the tag as it is updated. The cameras used in this deployment are separate from the PLC’s network. Images taken by the cameras can be acquired by connecting directly to them through a USB 3.0 wired connection. These cameras are industry-grade Basler cameras with a Python library, pyplon, which simplifies the image-capturing process. Since this is a wired connection, we can ensure that there is minimal communication lag time between the image request and capturing process. As such, once the cycle state tag is read as either four or nine, the images are captured from the cameras and sent into the model for detection.
Figure 6: Deployment Architecture
VIII Conclusion and Further Work
In this study, we curated an industry-standard dataset tailored for assembly processes using the novel YOLO-FF model and introduced AssemAI, a standardized image-based anomaly detection pipeline. We began with a simple CNN Model, which provided a baseline for anomaly detection through a straightforward architecture and standard training techniques. Building on this, we explored ViT and EfficientNet models to enhance classification accuracy and efficiency. A significant aspect of our approach was the focus on interpretability, where we aimed to understand both model and user lever explainability to high-level phenomena such as structural integrity. This not only improved model understanding but also provided actionable insights for anomaly detection. Our findings indicate that EfficientNet offers significant improvements over traditional methods. Future work should explore hybrid architectures, which take other modalities like time series and textual, further interpretability techniques and deployment on edge devices for real-time anomaly detection applications to extend these findings to other domains and production environments. Also, to improve how domain experts understand our approach, we suggest creating abstract representations of causal factors associated with anomalies. This approach focuses on linking sensor data to broader concepts like structural issues or gripper malfunctions, providing a more comprehensive view of anomalous events.
Acknowledgment
This work was supported in part by NSF grants #2119654, “RII Track 2 FEC: Enabling Factory to Factory (F2F) Networking for Future Manufacturing” and SCRA grant “Enabling Factory to Factory (F2F) Networking for Future Manufacturing across South Carolina”.

### v2

@@
 This paper introduces AssemAI, an interpretable image-based anomaly detection system tailored for smart manufacturing pipelines.
-Our primary contributions include the creation of a tailored image dataset and the development of a custom object detection model, YOLO-FF, designed explicitly for anomaly detection in manufacturing assembly environments.
-Utilizing the preprocessed image dataset derived from an industry-focused rocket assembly pipeline, we address the challenge of imbalanced image data and demonstrate the importance of image-based methods in anomaly detection.
+Utilizing a curated image dataset from an industry-focused rocket assembly pipeline, we address the challenge of imbalanced image data and demonstrate the importance of image-based methods in anomaly detection.
+Our primary contributions include deriving an image dataset, fine-tuning an object detection model YOLO-FF, and implementing a custom anomaly detection model for assembly pipelines.
 The proposed approach leverages domain knowledge in data preparation, model development and reasoning.
-We compare our method against several baselines, including simple CNN and custom Visual Transformer (ViT) models, showcasing the effectiveness of our custom data preparation and pretrained CNN integration.
-Additionally, we incorporate explainability techniques at both user and model levels, utilizing ontology for user-friendly explanations and SCORE-CAM for in-depth feature and model analysis.
-Finally, the model was also deployed in a real-time setting.
-Our results include ablation studies on the baselines, providing a comprehensive evaluation of the proposed system.
+We implement several anomaly detection models on the derived image dataset, including a Convolutional Neural Network, Vision Transformer (ViT), and pre-trained versions of these models.
+Additionally, we incorporate explainability techniques at both user and model levels, utilizing ontology for user-level explanations and SCORE-CAM for in-depth feature and model analysis.
+Finally, the best-performing anomaly detection model and YOLO-FF are deployed in a real-time setting.
+Our results include ablation studies on the baselines and a comprehensive evaluation of the proposed system.
 This work highlights the broader impact of advanced image-based anomaly detection in enhancing the reliability and efficiency of smart manufacturing processes.
+The image dataset, codes to reproduce the results and additional experiments are available at https://github.com/renjithk4/AssemAI.
 Index Terms:
-Anomaly Detection, Object Detection, Smart Manufacturing, Interpretability, Image Processing
+Smart Manufacturing, Object Detection, Image Processing, Anomaly Detection, Interpretability
 I Introduction
@@
 Anomaly detection is an essential methodology in manufacturing to identify deviations from normal production processes, which can indicate potential issues such as equipment failures, defects, or inefficiencies.
-Anomaly detection helps ensure product quality, reduce downtime and optimize production processes [3] [4].
+Anomaly detection helps ensure product quality, reduce downtime, and optimize production processes [3] [4].
 By monitoring various process parameters and product characteristics, manufacturers can detect anomalies early and prevent costly production disruptions [5].
-Image data is increasingly being leveraged in manufacturing systems due to advancements in computer vision and the availability of high-resolution cameras.
-In modern manufacturing facilities, image data is utilized for a wide range of applications, including quality inspection [6, 7], predictive maintenance [8, 9], process monitoring [10] and process optimization [11].
+Image data are increasingly being leveraged in manufacturing systems due to advancements in computer vision and the availability of high-resolution cameras.
+In modern manufacturing facilities, image data are utilized for a wide range of applications, including quality inspection [6, 7], predictive maintenance [8, 9], process monitoring [10] and process optimization [11].
 The integration of image data allows for real-time monitoring and anomaly detection, significantly improving the ability to identify defects and streamline production processes.
@@
 Automated visual inspection systems use image data to detect surface defects, assembly errors and deviations from design specifications[12], providing real-time feedback essential for maintaining high production standards and minimizing defects[4].
-Despite the advancements in image-based anomaly detection, several challenges persist.
-(i) High variability in appearance between normal and anomalous conditions, which can make it difficult for models to differentiate subtle defects from acceptable variations.
-(ii) Requirement for large and diverse datasets for training, which can be resource-intensive to collect and annotate [5, 13].
-(iii) Challenges with generalization, where a model trained on images from one manufacturing environment may not perform well in different settings due to variations in lighting, camera angles, or product types [14].
-(iv) Difficulty in interpreting detection results, affecting the ability to provide actionable insights for process improvement [15].
-Figure 1: Overall Architecture.
-The figure includes the pipeline for AssemAI, beginning with dataset preparation involving filtering and cropping images, followed by zero-shot object detection using two approaches (a custom model and YOLO-FF), and culminating in object detection modeling with CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET.
-The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations.
+Despite the advancements in image-based anomaly detection, several challenges persist as follows: (i) In visual data, the distinctions between normal and anomalous conditions are often subtle, making it difficult for the model to differentiate these anomalies from typical image variations (ii) The need for large and diverse datasets for training can be resource-intensive, both in terms of data collection and annotation [5, 13].
+(iii) Generalization issues that arise when a model trained in one manufacturing environment fails to adapt to others due to differences in lighting, camera angles, or product types [14].
+(iv) Difficulty in interpreting detection results, hindering the ability to provide actionable insights for process improvement [15].
 To address these challenges, this work presents the following contributions:
-  • A novel anomaly detection model utilizing the pre-trained EfficientNet architecture [16], fine-tuned to a new dataset for improved anomaly detection in assembly processes.
-  • A new image dataset for anomaly detection in manufacturing, curated using a custom-trained YOLOv9 model named YOLO-FF (You Only Look Once-Future Factories).
-  • Integration of advanced explainability techniques, incorporating an ontology-based method for user-level explanations and SCORE-CAM for detailed feature importance and model analysis.
-  • The image dataset and code to reproduce the results and additional experiments are available at this link.
+  1. 1.
+    A new image dataset for anomaly detection in manufacturing, derived using a zero-shot object detection model OWL-ViT.
+  2. 2.
+    A novel anomaly detection model that leverages the pre-trained EfficientNet architecture [16], fine-tuned on the derived dataset for enhanced anomaly detection in assembly pipelines.
+  3. 3.
+    Integration of user-level explainability using process ontology and model-level explainability using SCORE-CAM for salient feature and model analysis,
+Figure 1: Overall Architecture of AssemAI.
+The figure illustrates the AssemAI pipeline, beginning with dataset preparation, which includes filtering and cropping images.
+Next, object detection is performed using zero-shot detection and a fine-tuned model (YOLO-FF).
+This is followed by anomaly detection using CNN, Custom-ViT, Pre-trained-ViT, and EfficientNet.
+The detection output is explained using SCORE-CAM for model-level explanations and process ontology for user-level explanations.
 II Literature Review
 II-A Object Detection Models
-Object detection has evolved with various models, improving both accuracy and efficiency.
+Object detection has advanced significantly, with various models enhancing both accuracy and efficiency.
 Girshick et al.
@@
 [30] and Hayat et al.
-[31] advanced ZSD with DELO and Generative-ZSD.
+[31] advanced ZSD with Don’t Even Look Once (DELO) model and Generative-ZSD.
 Liu et al.
@@
 II-C Anomaly detection
-Anomaly detection has seen significant advancements across diverse domains like networking [34], smart agriculture [35], healthcare [36], foundational [35, 37, 38, 39], manufacturing [40, 41, 42, 43, 44, 45].
-Kim et al.
-introduced NetViewer [34], a network measurement approach that detects, identifies and visualizes attacks and anomalous traffic in real-time by passively monitoring packet headers and representing them as frames or images.
-An anomaly detection method using Generative Adversarial Networks (GAN), which searches for good representations of samples in the generator’s latent space and deems samples without such representations as anomalous, was proposed by Deecke et al.
-[37].
-Minhas et al, in their work addresses visual defect assessment for anomaly detection in tasks [38], like surface inspection by proposing a network-based deep transfer learning method using CNNs, hypothesizing that deeper networks will outperform traditional single class SVMs.
-II-D Anomaly detection in manufacturing
-Significant research has investigated image-based anomaly detection in manufacturing.
+Anomaly detection has seen significant advancements across diverse domains like networking [34], smart agriculture [35], healthcare [36], manufacturing [37, 38, 39, 40, 41, 42, 43].
 Haselmann et al.
@@
 The AssemAI pipeline begins with dataset preparation, where images are filtered and cropped to focus on relevant features.
-This is followed by zero-shot object detection using two approaches: a custom model and YOLO-FF, which adapts zero-shot detection for anomaly detecti use cases.
-Subsequently, object detection modeling is performed using several architectures, including CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET, to ensure robust anomaly detection.
-The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations, providing both technical and domain-specific insights.
-III-A Future Factories Multimodal Dataset
-We use the Future Factories (FF) dataset [46] generated and made publicly available by the Future Factories (FF) lab operating at the McNair Aerospace Research Center at the University of South Carolina.
-The dataset consists of measurements from a simulation of a rocket assembly pipeline, which adheres to industrial standards in deploying actuators, control mechanisms and transducers.
-The dataset is comprised of two versions; Analog and Multi-modal dataset, for which we use the images included in the Multi-modal dataset for our purpose.
-Table I shows the statistics of the original images in the FF Multimodal dataset.
+We then use a zero-shot object detection model to get the region of interest for the rocket in the image, deriving a dataset with masked regions of interest and bounding box details in a CSV file.
+A YOLO model is then fine-tuned on these images and bounding boxes.
+Subsequently, this derived dataset is used for anomaly detection experiments with several architectures, including CNN, Custom ViT, Pre-trained ViT, and EfficientNet.
 TABLE I: Artificats of the original images in FF Multimodal dataset
@@
 Types of classes 8
+The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations, providing technical and domain-specific insights.
+III-A Future Factories Multimodal Dataset
+We use the Future Factories (FF) dataset [45] curated and publicly released by the Future Factories lab at the McNair Aerospace Research Center, University of South Carolina.
+The dataset consists of measurements from a simulation of a rocket assembly pipeline, which adheres to industrial standards in deploying actuators, control mechanisms and transducers [^1] Refer to the images available in the GitHub repository for further details..
+The dataset comprises two versions; analog and multimodal dataset, for which we use the images included in the multimodal dataset in our study.
+Table I shows the statistics of the original images in the FF multimodal dataset.
+III-B Image filtering
+The rocket assembly process at the FF lab is divided into 21 distinct cycle states.
+Information about these cell cycle states is not directly available in the multimodal dataset and had to be extracted from the analog dataset using a mapping function provided by domain experts.
+By calculating the Structural Similarity Index (SSIM) between the normal and anomalous images as shown in Figure 2, we observed that the images are structurally very similar across most cycle states.
+This indicates that the differences between the images are subtle and localized, which suggests that further analysis is needed to filter and figure out the region of interest.
+Among these cell cycle states, the rocket and its parts are visible only in two specific cycle states due to the spatial location of robots and other machinery and the location of the cameras and camera angles.
+We focus on filtering cycle four and a section of cycle nine based on domain knowledge and observational insights for enhanced image understanding.
+Table II and III show the filtered dataset and its anomaly statistics, respectively.
+Figure 2: Structural Similarity between Normal and Anomalous Images
 TABLE II: Artificats of the filtered images in FF Multimodal dataset
@@
 count 12475 3119 15594
-III-B Image filtering
-The rocket assembly process at the FF lab is divided into 21 distinct cycle states.
-Information about these cell cycle states was not directly available in the multi-modal dataset and had to be extracted from the analog dataset using a mapping function provided by domain experts.
-By calculating the Structural Similarity Index (SSIM) between the normal and anomalous images as shown in Figure 2, we observed that the images are structurally very similar across most cycle states.
-This indicates that the differences between the images are subtle and localized, which suggests that further analysis is needed to filter and figure out the region of interest.
-Among these cell cycle states, the rocket and its parts are visible only in two specific cycle states due to the spatial location of robots and other machinery and the location of the cameras and camera angles.
-We focus on filtering cycle four and a section of cycle nine based on domain knowledge and observational insights for enhanced image understanding.
-Table II and III show the filtered dataset and its anomaly statistics, respectively.
-Figure 2: Structural Similarity between Normal and Anomalous Images
 III-C Image Cropping
 Our dataset comprises images captured at various stages of the manufacturing cycle.
-Each cycle consists of defined time stamps where the toy rocket’s position remains consistent.
-Leveraging this domain knowledge, we have implemented a cropping strategy to remove the background from all images.
-This consistent positioning of the toy rockets across cycles enables us to crop the images accurately, ensuring that only the relevant parts of the images are retained.
-The necessity for this cropping approach became evident after our initial attempts at anomaly detection without cropping yielded sub-optimal results.
+In each cycle, the rocket’s position remains consistent within each respective state.
+Using this domain knowledge, we defined a cropping strategy by defining a bounding box for each cycle state to remove the background from all images.
+This consistent positioning of the rockets across cycles enables us to crop the images accurately, ensuring that only the relevant parts of the images are retained.
+The suboptimal results obtained with uncropped images necessitated cropping of images.
 Specifically, our models performed poorly, with significant misclassification rates.
 We employed the explainability model SCORE-CAM [46] to understand the underlying issues.
-The insights provided by this model revealed that the models disproportionately focused on the images’ background rather than on the toy rockets themselves, as shown in Figure 3.
-This misdirection was the primary cause of the poor performance.
-By cropping the images to remove the background, we ensured that the models concentrated on the relevant features of the toy rockets.
-This adjustment improved the model performance, leading to more accurate anomaly detection.
-The final cropped images are of two sizes, determined by the cycle states: 200 pixels by 70 pixels for cycle state 4 and 400 pixels by 205 pixels for cycle state 9.
-Figure 3: Score-CAM visualization highlighting the model’s focus on background elements emphasizing the need for cropping and object detection to improve accuracy by isolating the toy rocket.
+The insights provided by SCORE-CAM revealed that the models prioritized images’ backgrounds rather than the rockets themselves, as shown in Figure 3.
+Therefore, images were cropped to remove the backgroun.
+This image transformation improved the model performance, leading to better performance.
+As the rocket is the most important feature, the images are cropped accordingly, which resulted in a 200x70 pixel image for cycle state four and 400x205 pixels for cycle state nine.
+Figure 3: Score-CAM visualization shows the model mistakenly focusing on background elements, highlighted in red as the most important regions, rather than the assembly pipeline marked in blue.
+This emphasizes the need for cropping and object detection to improve accuracy by isolating the relevant parts of the image.
 III-D Text Guided Zero-Shot Object Detection
-We then incorporated a zero-shot object detection model to enhance the accuracy and robustness of our anomaly detection pipeline targeting the detection of toy rockets.
-Specifically, we employed OWL-ViT [48], a model that leverages multi-modal representations to perform open-vocabulary detection.
+We then incorporate a zero-shot object detection model to enhance the accuracy and robustness of our anomaly detection pipeline, which targets the detection of rockets and their parts.
+Specifically, we employ OWL-ViT [47], a model that leverages multimodal representations to perform open-vocabulary detection.
 This approach allows for the detection of objects based on free-text prompts, facilitating a flexible and powerful detection mechanism.
 OWL-ViT integrates CLIP (Contrastive Language-Image Pretraining) with lightweight object classification and localization heads.
-This integration enables the model to handle open-vocabulary detection by embedding free-text queries through CLIP’s text encoder, which are then utilized as inputs for the object classification and localization heads.
-The Vision Transformer (ViT) processes image patches as inputs, associating them with their corresponding textual descriptions.
-III-E Training a Custom Object Detection Model with YOLOv9
-To further enhance our anomaly detection pipeline for toy rockets, we trained a custom object detection model using YOLOv9 [19].
-While zero-shot models like OWL-ViT offer significant flexibility and adaptability, a custom-trained model provides specific advantages such as:
-  • By training YOLOv9 specifically on our dataset of toy rockets, we can fine-tune the model parameters to achieve higher accuracy and precision for this specific task, compared to a more general-purpose zero-shot model.
-  • YOLOv9 is designed for high-speed inference, making it ideal for real-time applications in a manufacturing pipeline.
-In contrast, zero-shot models like OWL-ViT, while flexible, may have higher computational requirements and longer inference times.
+This integration enables the model to handle open-vocabulary detection by embedding free-text queries through CLIP’s text encoder.
+These are then utilized as inputs for the object classification and localization heads.
+The ViT processes image patches as inputs, associating them with their corresponding textual descriptions.
+III-E Fine-tuning Object Detection Model with YOLOv9
+To further enhance the object detection pipeline for rocket assembly, we fine-tune a YOLOv9 object detection model [19] using the dataset derived from OWL-ViT, which we refer to as YOLO-FF.
+While zero-shot models like OWL-ViT offer significant flexibility and adaptability, a fine-tuned model provides specific advantages as follows:
+  • Fine-tuning YOLOv9 using the derived dataset ensures the model parameters are optimized for the specific task, facilitating better accuracy and precision.
+  • YOLOv9 is optimized for high-speed inference, making it ideal for real-time applications, whereas models like OWL-ViT typically require more computation and have longer inference times.
 IV Problem Formulation
-Consider a dataset $\mathcal{D}$ comprising images captured at various stages of the toy rocket manufacturing process.
+Consider a dataset $\mathcal{D}$ comprising images captured at various stages of the rocket assembly process.
 The manufacturing process is divided into cycles, each consisting of 21 distinct states.
-Let $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$ denote the set of manufacturing cycles, where each cycle $C_{j}$ for $j=1,2,\ldots,n$ represents the complete assembly of one toy rocket.
+Let $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$ denote the set of manufacturing cycles, where each cycle $C_{j}$ for $j=1,2,\ldots,n$ represents the complete assembly of one rocket.
 Each cycle $C_{j}$ is divided into 21 states, represented by the set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{21}\}$.
@@
 Additionally, let $B_{j,s}=(x_{\min},y_{\min},x_{\max},y_{\max})$ denote the bounding box for image $I_{j,s}$, obtained via object detection techniques to focus on the relevant part of the image.
-We define the dataset as:
-$\mathcal{D}=\{(I_{j,s},y_{j,s},B_{j,s})\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$
+We define the dataset as in equation 1.
+$\mathcal{D}=\{(I_{j,s},y_{j,s},B_{j,s})\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$ (1)
 IV-A Task Description
@@
   2. 2.
-    Cropping each image $I_{j,s}$ using its bounding box $B_{j,s}$ which is mapped to its specific state. This effectively isolates the regions of interest while removing background noise.
+    Cropping each image $I_{j,s}$ using its bounding box $B_{j,s}$, which is mapped to its specific state. This effectively isolates the regions of interest while removing background noise.
   3. 3.
-    Applying AssemAI to classify the cropped images.
-The primary objective is to detect anomalies in the images, leveraging the bounding boxes to enhance the model’s ability to focus on the critical parts of the image.
+    Detecting objects in the cropped images using an object detection model. This will generate the region of interest (i.e., rocket parts) and their bounding boxes.
+  4. 4.
+    Classifying anomalies from the detected objects using an anomaly detection model.
+The main goal is to detect anomalies in the images by using bounding boxes to help the model focus on the important parts.
 To address class imbalance, we use a weighted cross-entropy loss, where higher weights are assigned to classes with fewer images.
-Mathematically, we aim to minimize the weighted classification error defined as:
-$\min_{\theta}\frac{1}{n}\sum_{j=1}^{n}\sum_{s\in\{4,9\}}w_{y_{j,s}}\mathcal{L}(f(I_{j,s};\theta),y_{j,s})$
-where $\theta$ represents the model parameters, $\mathcal{L}$ is the cross-entropy loss function and $w_{y_{j,s}}$ is the weight assigned to the class $y_{j,s}$.
+We aim to minimize the weighted classification error defined in equation 2:
+$\min_{\theta}\frac{1}{n}\sum_{j=1}^{n}\sum_{s\in\{4,9\}}w_{y_{j,s}}\mathcal{CE}(f(I_{j,s};\theta),y_{j,s})$ (2)
+where $\theta$ represents the model parameters, $\mathcal{CE}$ is the cross-entropy loss function, and $w_{y_{j,s}}$ is the weight assigned to the class $y_{j,s}$.
 By focusing on specific cycle states and using object detection to preprocess the images, we aim to improve the accuracy and reliability of our anomaly detection system.
-The ultimate goal is to create a robust model that accurately detects anomalies in the toy rocket manufacturing process.
 V Experiments
 We outline the following experimental setup to evaluate the AssemAI’s overall performance and the contribution of each sub-module.
-V-A Common Hyperparameters and Training Setup
-The preprocessed dataset was split into training (80%) and testing (20%) sets.
-For all models, we used Cross-Entropy Loss and the Adam optimizer [49], tuning hyperparameters like epochs, batch size and learning rate.
-The best model was saved based on validation accuracy.
-The dataset preprocessing involved resizing images to the required input dimensions and normalizing them according to the specific requirements of each model.
-V-B Baselines
-V-B1 Custom CNN
-The Simple CNN model architecture included two convolutional layers (32 and 64 filters respectively), each followed by a max-pooling layer.
-After flattening, the output was passed through a fully connected layer with 512 neurons, then to the final layer corresponding to the number of classes (5).
-V-B2 Custom ViT
-The Vision Transformer (ViT) model [50] was implemented from scratch, starting with a patch embedding layer that splits the input image into non-overlapping patches, each embedded into a high-dimensional space.
-This was followed by a series of transformer blocks for learning spatial relationships within the image patches.
-Each transformer block consisted of a multi-head self-attention mechanism and a feed-forward neural network, with layer normalization and residual connections applied at each step.
-The final classification head mapped the output to the desired number of classes (5).
-V-B3 Pretrained ViT
-We utilized the pre-trained Vision Transformer (ViT) model, specifically google/vit-base-patch16-224.
-Images were pre-processed by dividing each image into a sequence of fixed-size non-overlapping patches, then linearly embedded.
-A [CLS] token was added to represent the entire image, facilitating classification.
-Absolute position embeddings were incorporated and the resulting sequence of vectors was fed to the standard Transformer encoder.
-The ViTImageProcessor resized and normalized images to the required 224x224 resolution.
-The model was loaded with half-precision (torch.float16).
-Training utilized PyTorch Lightning’s Trainer, with a weighted Cross-Entropy loss function and the AdamW optimizer with a learning rate scheduler.
-The best checkpoint was saved by monitoring the validation loss.
-V-C Proposed Approach
-We implemented an EfficientNet-B0 model [16] for image classification tasks.
-The EfficientNet-B0 architecture was selected for its state-of-the-art performance and efficiency in image classification.
+V-A Experiments with YOLO-FF Model
+The YOLOv9 model is fine-tuned on the derived labeled image dataset resulting from the OWL-ViT model.
+It is trained for 50 epochs.
+During training, the model achieved an accuracy of approximately 99% on the validation dataset.
+V-B Common Hyperparameters and Training Setup
+The derived image dataset is split into training (80%) and testing (20%) sets.
+We used Cross-Entropy Loss and the Adam optimizer [48] for baselines and the proposed models, tuning hyperparameters like epochs, batch size and learning rate.
+The best model is saved based on validation accuracy.
+The dataset preprocessing involved resizing images to the required input dimensions (224×224 pixels), normalizing ([0.485,0.456,0.406]) and standardizing ([0.229,0.224,0.225]) according to the specific requirements of each model.
+V-C Baselines
+V-C1 Custom CNN
+The Simple CNN model architecture includes two convolutional layers (32 and 64 filters respectively), each followed by a max-pooling layer.
+After flattening, the output is passed through a fully connected layer with 512 neurons, then to the final layer corresponding to the number of classes (5).
+V-C2 Custom ViT
+The Vision Transformer (ViT) model [49] is implemented from scratch, starting with a patch embedding layer of size 16 that splits the input image into non-overlapping patches, each embedded into a high-dimensional space of size 768.
+This is followed by a series of transformer blocks for learning spatial relationships within the image patches.
+Each transformer block consists of a multi-head self-attention mechanism and a feed-forward neural network, with layer normalization and residual connections applied at each step.
+The final classification head maps the output to the desired number of classes (5).
+V-C3 Pre-trained ViT
+We utilize the pre-trained Vision Transformer (ViT) model, specifically google/vit-base-patch16-224.
+Images are preprocessed by dividing each image into a sequence of fixed-size non-overlapping patches, then linearly embedded.
+A [CLS] token is added to represent the entire image, facilitating classification.
+Absolute position embeddings are incorporated and the resulting sequence of vectors is fed to the standard Transformer encoder.
+The ViTImageProcessor resizes and normalizes images to the required 224x224 resolution.
+The model is loaded with half-precision (torch.float16).
+The training utilized PyTorch Lightning’s Trainer, with a weighted Cross-Entropy loss function, and the AdamW optimizer with a learning rate scheduler.
+The best checkpoint is saved by monitoring the validation loss.
+Figure 4: Experimental results with OWL-ViT. Class 1-5 denotes anomaly types: [No Anomaly], [NoNose], [NoNose,NoBody2], [NoNose,NoBody2,NoBody1] and [NoBody1] respectively
+V-D Proposed Modeling Approach
+We implement an EfficientNet-B0 model [16] for image classification tasks.
+The EfficientNet-B0 architecture is selected for its state-of-the-art performance and efficiency in image classification.
 It leverages a compound scaling method to optimize model depth, width and resolution, achieving high accuracy with fewer parameters compared to traditional networks.
-The model was pre-trained on ImageNet and adapted for our task by modifying the final classification layer to match the number of classes in our dataset (5 classes).
-The model was trained for 10 epochs, with the best model saved based on validation accuracy.
+The model is pre-trained on ImageNet and adapted for our task by modifying the final classification layer to match the number of classes in our dataset (5 classes).
 VI Results
@@
 The EfficientNet model achieves 96% overall accuracy and 95% of precision, 96% of recall, 95% of F1 score, respectively.
-Figure 4 depicts the performance of detecting various anomaly types and the normal class across various modeling approaches.
-It can be observed that among all the models, EfficientNet with the YOLO-FF module gives the best results in detecting all the types of classes.
-TABLE IV: Experimental Results
-OD Model
-WP
-WR
-WF1
-Accuracy Support*
-Simple CNN
-+YOLO-FF
-92.00±
+Figure 4 depicts the performance of detecting various anomaly types and the normal class across various modeling approaches with the YOLO-FF object detection approach.
+It can be observed that among all the models, EfficientNet gives the best results in detecting all the types of classes.
+TABLE IV: Experimental Results of AssemAI
+Model WP WR WF1 Accuracy Support*
+Simple CNN*
+82.00±
 1.00%
-91.00±
+81.00±
 1.00%
-91.00±
+83.00±
 1.00%
-91.00±
+82.05±
 1.00%
 3119
-Custom ViT
-+YOLO-FF
-71.00±
+Custom ViT*
+85.00±
 1.00%
-66.00±
+86.00±
+0.05%
+83.00±
 1.00%
-66.00±
-1.00%
-65.00±
+84.10±
 1.00%
 3119
-Pretrained ViT
-+YOLO-FF
+Pre-trained ViT*
 88.50±
@@
 3119
-EfficientNet
-+YOLO-FF
+EfficientNet*
 95.00±
@@
 3119
-*Support-Number of samples used. WP-Weighted Average Precision, WR-Weighted Average Recall, WF1-Weighted Average F1-Score
-Figure 4: Experimental results. Class 1-5 denotes anomaly types: [No Anomaly], [NoNose], [NoNose,NoBody2], [NoNose,NoBody2,NoBody1] and [NoBody1] respectively
+Models marked in * are experimented with the derived image dataset.
+EfficientNet model is experimented for original images (includes all cycle states) and for filtered images (includes cycle states four and nine).
+Bold indicates the best performance.
 VI-A Additional experiments on Explainability
 VI-A1 User level explainability
-In this work, we employ process ontology designed and developed for Future Factories Rocket assembly line.
-In comparison with conventional ontologies, process ontology not only captures the definition of sensors but also captures the procedural nature of the assembly process.
-This aids in understanding the involvement of sensors and equipments at a given point in assembly process.
-The Future Factories assembly is divided into 21 cycle states which forms the basis of process ontology construction.
-The specific features of process ontology are as follows: (i) consists of definition and item specification of sensors and equipments, (ii) relationship between the sensors and equipments, (iii) function and involvement of each sensors and robots with respect to the cycle states (iv) expected (or anomalous) values of sensor variables in with respect to each cycle state (v) type of anomaly that could be associated with each cycle state (iv) sensor values and other knowledge can be dynamically updated as per the change in experiment set up.
+In this work, we employ process ontology designed and developed for the Future Factories Rocket assembly line.
+In contrast to conventional ontologies, process ontology not only captures the definition of sensors but also captures the procedural nature of the assembly process.
+This aids in understanding the involvement of sensors and types of equipment at a given point in the assembly process.
+The Future Factories assembly is divided into 21 cycle states, which form the basis of the ontology construction process.
+The specific features of process ontology are as follows: (i) consists of definition and item specification of sensors and types of equipment, (ii) relationship between the sensors and types of equipment, (iii) function and involvement of each sensor and robot with respect to the cycle states (iv) expected (or anomalous) values of sensor variables in with respect to each cycle state (v) type of anomaly that could be associated with each cycle state (iv) sensor values and other knowledge can be dynamically updated as per the change in experiment set up.
 Capturing the procedural nature of the assembly process aids in understanding the contribution of sensors in anomalies.
-For example, if the initial stage of assembly line is being analyzed for anomaly, it can be understood from ontology that Robot-4 and its corresponding sensors do not contribute to this anomaly.
+For example, if the initial stage of the assembly line is being analyzed for anomaly, it can be understood from the ontology that Robot-4 and its corresponding sensors do not contribute to this anomaly.
 The goal of process ontology is to explain and assess the output from the models.
@@
 On the other hand, process ontology can also verify if the model predictions are correct in certain cases (Figure 5).
-Since each image is associated with a timestamp which in turn can be mapped to cycle state, the predicted output from the model can be verified using anomaly types defined in the process ontology.
-Illustrating using an output from the model using Figure 5, we notice that the input image is associated with cycle state 4.
+Since each image is associated with a timestamp which in turn can be mapped to the cycle state, the predicted output from the model can be verified using anomaly types defined in the process ontology.
+Figure 5 illustrates that the input image is associated with cycle state 4.
 The model predicted the image to be anomalous with type NoNose.
-This is an incorrect prediction as per the ontology as NoNose anomaly can happen only from cycle state 8 onwards.
-Using this knowledge and verifying the outputs model predictions, it is found that the model incorrectly predicted NoBody2 anomaly 105 out of 801 times and NoNose anomaly 245 out of 1145 times.
+This is an incorrect prediction as per the ontology as NoNose anomaly can happen only from cycle stage 8 onwards.
+Using this knowledge and verifying the outputs of model predictions, it is found that the model incorrectly predicted NoBody2 anomaly 51 out of 801 times and NoNose anomaly 106 out of 1145 times.
+The ontology also acts as an additional layer that catches the misclassification of the model, enhancing the robustness of the proposed anomaly detection pipeline.
 Figure 5: Verification of the explanations through process ontology
+Figure 6: Deployment Setup of AssemAI
+Figure 7: Web User interface of AssemAI
 VII Deployment of AssemAI
-After training and testing the model, we then worked on deploying AssemAI on the Future Factories lab.
-This deployment strategy can be seen in 6.
-AssemAI requires two separate types of inputs to operate in real-time.
+After training and testing the model, we deploy AssemAI on the Future Factories lab.
+This deployment setup can be seen in Figure 6.
+AssemAI requires two separate types of inputs to operate in real time.
 The script that utilizes the model should be able to obtain the current cycle state of the assembly process.
-This data tag allows the script to ensure that the images captured and input into the model are from cycle states 4 and 9.
+This data tag allows the script to ensure that the images captured and input into the model are from cycle states four and nine.
 This tag is obtained by connecting to the OPC UA server running on the Programmable Logic Controller (PLC) and constantly reading the tag as it is updated.
@@
 These cameras are industry-grade Basler cameras with a Python library, pyplon, which simplifies the image-capturing process.
-Since this is a wired connection, we can ensure that there is minimal communication lag time between the image request and capturing process.
+Since this is a wired connection, we can ensure that there is minimal communication lag time between the image request and the capturing process.
 As such, once the cycle state tag is read as either four or nine, the images are captured from the cameras and sent into the model for detection.
-Figure 6: Deployment Architecture
+Figure 7 illustrates the web user interface of the deployed AssemAI.
 VIII Conclusion and Further Work
-In this study, we curated an industry-standard dataset tailored for assembly processes using the novel YOLO-FF model and introduced AssemAI, a standardized image-based anomaly detection pipeline.
+In this study, we derived an industry-standard dataset tailored for assembly processes using the novel YOLO-FF model and introduced AssemAI, a standardized image-based anomaly detection pipeline.
 We began with a simple CNN Model, which provided a baseline for anomaly detection through a straightforward architecture and standard training techniques.
 Building on this, we explored ViT and EfficientNet models to enhance classification accuracy and efficiency.
-A significant aspect of our approach was the focus on interpretability, where we aimed to understand both model and user lever explainability to high-level phenomena such as structural integrity.
+A significant aspect of our approach is the focus on interpretability, where we aimed to understand both model and user lever explainability to high-level phenomena such as structural integrity.
 This not only improved model understanding but also provided actionable insights for anomaly detection.
@@
 Acknowledgment
-This work was supported in part by NSF grants #2119654, “RII Track 2 FEC: Enabling Factory to Factory (F2F) Networking for Future Manufacturing” and SCRA grant “Enabling Factory to Factory (F2F) Networking for Future Manufacturing across South Carolina”.
+This work is supported in part by NSF grants #2119654, “RII Track 2 FEC: Enabling Factory to Factory (F2F) Networking for Future Manufacturing” and SCRA grant “Enabling Factory to Factory (F2F) Networking for Future Manufacturing across South Carolina”.